OpenStudy (anonymous):

Find the distance from P to l. Line l contains points (6,5) and (2,3). Point P has coordinates (2,6).

OpenStudy (amistre64):

vector from the line, and vector from a line point to the other point will be useful

OpenStudy (amistre64):

(6,5) and (2,3). (2,6) -2-3 -2-3 -2-3 -------------------- 4,2 0,0 0,3 and the length of each vector sqrt(20) 3 \[cos(a)=\frac{u.v}{|u|~|v|}\] \[cos(a)=\frac{<4,2>\cdot <0,3>}{6 \sqrt5}\] \[cos(a)=\frac{6}{6 \sqrt5}\] \[a=cos^{-1}\left(\frac1{\sqrt5} \right) \]

OpenStudy (amistre64):

sin of a times length of point vector should equal distance from line to point

OpenStudy (amistre64):



OpenStudy (amistre64):

another way would be to define the slope of the line, and perp it to get the line of the outer point; then see where the 2 lines cross and define the distance between those points

OpenStudy (amistre64):

which method looks more familiar to you?